OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, 6) and has a slope of –1?

OpenStudy (anonymous):

You should know how to do this by now. Try to understand it. y = mx + b M is always slope. y = -1x + b Plug in your points x, y into this. 6 = -1(-2) + b Solve for b. b = 4 Plug b back into y = -1x +b. Answer is y = -x + 4

OpenStudy (anonymous):

If you want ax + by = c form, you move -x over to the other side. x + y = 4